when three numbers are added in pairs,the sums of the pairs are,22,39,and 45. What are the three numbers?
Answer by ankor@dixie-net.com(22740)

when three numbers are added in pairs,the sums of the pairs are,22,39,and 45. What are the three numbers?
:
a+b = 22
a+c = 39
b+c = 45
:
Arrange the last two equations for elimination
b + c = 45
a + c = 39
-------------subtraction eliminates c
b - a = 6
b = a+6
In the first equation, replace b with a+6
a + (a+6) = 22
2a = 22 - 6
2a = 16
a = 8
Use the 2nd equation to find c,
8 + c = 39
c = 39-8
c = 31
Use the last equation to find b
b + 31 = 45
b = 45-31
b = 14
:
We have: a=8, b=14, c=31